A3B2 is a sparingly soluble salt of molar mass M(g/mol) and solubility x g/litre. The ratio of the molar concentration of B3- to the solubility product of the salt is?
 

Grade:10

1 Answers

Arun
5 years ago
Dear student,
A3B2------> 3A+2 + 2B-3
Let the solubility is x g/L
Then,
Ksp = (3s)3(2s)2 = 108 x5
Ionic product = 108 (x5/M5)

Ionic product/ molar concentration of B-3 = 
 
108/2 x⁴/M⁴ = 54 x⁴/M⁴